# Scalar-induced gravitational waves including isocurvature perturbations with lattice simulations

### v1

@@ 5 Discussion and conclusions
The inclusion of the energy transfer term enables a more realistic treatment of the matter-radiation system in eMD scenarios. We find that a larger initial PBH abundance or mass results in a steeper power-law spectrum near the peak. In contrast, for non-topological solitons (Q-balls), a faster decay rate (corresponding to a smaller $n$ in Eq. (4.13)) yields a higher amplitude and a steeper slope near the peak. While the present study considers PBHs with very small masses, our results are directly relevant for guiding future analytical and numerical work. It is important to stress that our lattice simulation solves the perturbative equations of motion. Therefore, the present analysis does not account for phenomena like the nonlinear cutoff scale [52] or dynamics that lie outside the perturbative regime. Addressing this regime likely requires full numerical-relativity simulations or, at a minimum, the inclusion of the acoustic gravitational waves discussed in Refs. [70, 71].

### v2

@@ 5 Discussion and conclusions
The inclusion of the energy transfer term enables a more realistic treatment of the matter-radiation system in eMD scenarios. We find that a larger initial PBH abundance or mass results in a steeper power-law spectrum near the peak. In contrast, for non-topological solitons (Q-balls), a faster decay rate (corresponding to a smaller $n$ in Eq. (4.13)) yields a higher amplitude and a steeper slope near the peak. While the present study considers PBHs with very small masses, our results are directly relevant for guiding future analytical and numerical work. It is important to stress that our lattice simulation solves the perturbative equations of motion. Therefore, the present analysis does not account for phenomena like the nonlinear cutoff scale [9] or dynamics that lie outside the perturbative regime. Addressing this regime likely requires full numerical-relativity simulations or, at a minimum, the inclusion of the acoustic gravitational waves discussed in Refs. [82, 62, 61].
@@ Appendix A Semi-analytical formulas in isocurvature case
Appendix B The effect of relative velocity
In this section, we tested the effect of the relative velocity. We observe that the effects are more prominent in the early stage of the scenario but are quickly overwhelmed by other contributions in the later stage, as shown in Fig. 8.
Figure 8: The GW energy spectra at different times. Left: $M_{\mathrm{PBH}}=10^{-4}g$, $\beta=10^{-2}$. Right: $M_{\mathrm{PBH}}=2\times 10^{-4}g$, $\beta=10^{-2}$. We have also tested different $\beta$, and they show similar behaviors. The solid lines denote the results with relative velocity terms, while the dashed lines describe the results without relative velocity terms.